A monopolistically competitive firm faces the following demand curve for its product: 6 Price ($) Quantity 10 2 9 4 8 6 7 8 5A monopolistically competitive firm faces the following demand curve for its product: 6 Price ($) Quantity 10 2 9 4 8 6 7 8 5 12 4 14 3 16 2 18 1 20 10 Refer to the Table. The firm has total fixed costs of $20 and a constant marginal cost of $5 per unit. What will the firm do? a) It will produce 2 units; firms will exit the market in the long run. b) It will produce 4 units: firms will exit the market in the long run. *) It will produce 6 units; firms will exit the market in the long run. d) It will produce 8 units: firms will exit the market in the long run.
Answer #1

Answer d. it will produce 6 units and the firm will exit the long run.

Reason- MC=MR when Q=6

When Q=6

TR= 8*6=48

When Q=4

TR= p*Q= 4*9=36

MR when Q=6 = ∆TR/∆Q= (48-36)/(6-4)= 12/2=Since MR≈MC=5 when Q=6, optimal level is when Q6.

When Q=6, Profit= TR-TC= 48- 20-(5*6)= -$2

So there is a loss of $2. So firms will exit the market in the long run.